# This manuscript is a preprint and has not been peer reviewed | Version: July 2025

### v1

@@ Abstract
Deep Neural Networks often suffer from a critical limitation known as Catastrophic Forgetting, where performance on past tasks degrades after learning new ones. This paper introduces a novel continual learning approach inspired by human learning strategies like Active Recall, Deliberate Practice and Spaced Repetition, named Task-Focused Consolidation with Spaced Recall (TFC-SR). TFC-SR enhances the standard experience replay with a mechanism we termed the Active Recall Probe. It is a periodic, task-aware evaluation of the model’s memory that stabilizes the representations of past knowledge. We test TFC-SR on the Split MNIST and the Split CIFAR-100 benchmarks against leading regularization-based and replay-based baselines. Our results show that TFC-SR performs significantly better than these methods. For instance, on the Split CIFAR-100, it achieves a final accuracy of 13.17% compared to standard replay’s 7.40%. We demonstrate that this advantage comes from the stabilizing effect of the probe itself, and not from the difference in replay volume. Additionally, we analyze the trade-off between memory size and performance and show that while TFC-SR performs better in memory-constrained environments, higher replay volume is still more effective when available memory is abundant. We conclude that TFC-SR is a robust and efficient approach, highlighting the importance of integrating active memory retrieval mechanisms into continual learning systems.
Keywords Continual Learning  $\cdot$ Catastrophic Forgetting  $\cdot$ Deep Neural Networks  $\cdot$ Replay-based Continual Learning
@@ 1 Introduction
Deep Neural Networks often suffer from a critical limitation known as Catastrophic Forgetting: performance on previously learnt tasks degrades as the model learns to perform newer ones [1]. This problem arises from the Stability-Plasticity dilemma [2] where a model must be flexible enough to learn newer tasks while simultaneously being stable enough to retain existing knowledge. This seriously affects the ability of a model to perform well in real life scenarios where the model must adapt to newer data and tasks incrementally over a period of time.
This paper aims to test an approach inspired by effective, synergistic learning strategies employed by humans, such as Active Recall [3], Deliberate Practice [4], and Spaced Repetition [5]. Active Recall involves effortfully retrieving information from memory. Deliberate practice involves studying a task until a certain level of proficiency is achieved. Spaced repetition refers to reviewing information at increasing intervals with the interval length determined by how well the task has been mastered. For example, If a student is learning Japanese, they may use flashcard applications like Anki to frequently test harder vocabulary (Active Recall), while words that are easier and recalled frequently are studied at progressively longer intervals (Spaced Repetition).
This paper proposes and evaluates a novel methodology to replicate these human learning methods within the continual learning framework. The primary contribution is Task-Focused Consolidation with Spaced Recall (TFC-SR), where the model learns new tasks while concurrently “practicing” previously learnt tasks using a form of experience replay.The core of our method is the Active Recall Probe, a periodic ’memory check’ where the model performs a forward pass on past experiences to evaluate its own memory state. The outcome of this memory check is then used by an adaptive spaced schedule that determines the intensity and frequency of future checks. For this initial investigation, A “naive” notion of mastery is used where mastery refers to retained performance on previously learnt tasks exceeding a predefined threshold. For the purpose of validation, TFC-SR is tested on the Split MNIST and then on the more challenging Split CIFAR-100 benchmarks and compared against standard continual learning baselines, including premier regularization-based methods like EWC [6] and SI [7] to demonstrate its effectiveness in alleviating Catastrophic Forgetting.
@@ 2.1 Task Protocols
For the purpose of demonstrating the effectiveness of TFC-SR, a simple CNN and a Resnet-18 [8] were tasked to work on two benchmark datasets, Split MNIST [9] and Split CIFAR-100 [10], respectively. In both cases, after the model had learnt a task, it had to classify all the classes it had seen so far. This was done to test how well the model retained its old knowledge after new information has been made available to it.
Training a simple CNN on MNIST is a relatively simple task and can often be described as being the “hello, world!” of Neural Networks. A basic implementation can be easily trained to achieve high accuracy on the dataset and thus, It was used to test the feasibility of TFC-SR in a relatively simple setting. MNIST consists of a total of 70000 grayscale, 28x28 images of handwritten digits 0-9. Those 70000 images are then split into two sets - 60000 for training and 10000 for testing. Split MNIST is a version of this dataset that is further divided into multiple classification tasks. In our case, It was divided into 5 tasks where each task involved the model learning and classifying a pair of digits. This was done to demonstrate continual learning where the model gets data sequentially instead of having all the training data available at once.
CIFAR-100 was chosen to represent a better, relatively more difficult benchmark after TFC-SR demonstrated effectiveness within a simple setting. CIFAR-100 consists of 60000 color, 32x32 images representing 100 classes. There are 600 images per class and the 100 classes are grouped into 20 different superclasses. Similar to MNIST, CIFAR-100 is also split into two sets of training and testing sets with each containing 50000 and 10000 images respectively. Since this was going to be too difficult for a simple CNN to handle, a Resnet-18 was chosen as the model of choice. For our purposes, CIFAR-100 was divided into 10 tasks, with 10 classes per task.
@@ 2.2 Model Architecture
All models were implemented using Pytorch and were trained from scratch. For the first benchmark of Split MNIST, a standard CNN was used with convolution blocks followed by a classifier head. The first convolution block consisted of a 2D convolution layer with 32 3x3 kernels followed by a ReLU activation function and a 2x2 Max Pooling layer. The second convolution block was identical to the first one with the only difference being that the 2D convolution layer had 64 3x3 kernels. The output from these blocks would then be flattened and then passed to the classifier head, which consisted of two fully connected layers. The first fully connected layer 128 units with ReLU activation while the second one was a linear layer with 10 units representing the 10 digit classes.
For Split CIFAR-100, a ResNet-18 model was utilized. To adapt the model to the CIFAR dataset’s 32x32 image size, two conventional modifications were made: The first being that the initial 7x7 convolutional layer was replaced with a 3x3 convolution layer, and the subsequent max-pooling layer was removed by replacing it with an identity mapping. The final fully connected layer was replaced with a new linear layer with 100 units to represent the 100 classes in the CIFAR-100 dataset.
@@ 2.3 Learning Algorithms and Baselines
  • Sequential Fine tuning (Baseline). This represented the base, low bound benchmark without any additional method and was used to evaluate if there was any benefit to using TFC-SR at all.
  • Standard Experienced Replay [11] where the model had access to data from past tasks. Since TFC-SR is also a form of replay based method, Standard ER helped provide a better view into how the addition of Active Memory Probe helps.
  • Elastic Weight Consolidation (EWC) [6] is a premier regularization based method that adds a regularization loss in order to protect weights that are important for previous tasks based on Fisher Information Matrix.
  • Synaptic Intelligence (SI) [7] is another leading regularization based method. SI protects weights that have contributed significantly to reducing the loss function during training of past tasks. The regularization strength, $\lambda$, was found using coarse hyperparameter tuning for both EWC and SI.
@@ 2.4 TFC-SR algorithm
TFC-SR enhances the standard experience replay framework using an adaptive, scheduling mechanism based on active recall and spaced repetition. The training process for each task consists of two main components: continuous mixed batch training and an adaptive active recall schedule.
Continuous Mixed Batch training: For all tasks, the model is trained on mixed batches of data. Each batch is composed of a 50/50 split of new examples from the current task and examples randomly sampled from a replay buffer containing data from all the past tasks. To maintain a diverse and representative memory of the past experiences within a fixed memory budget, a Reservoir Replay Buffer of a fixed capacity was used.
Active Recall Probe and Adaptive Scheduling: The core novelty of TFC-SR is its adaptive scheduling which we term Active Recall Probe. This probe is a “memory check” performed periodically after a training epoch. It consists of evaluating the model’s performance on a batch of samples from the replay buffer. To get an accurate measure of retained knowledge in a class-incremental setting, this evaluation is performed in a task-aware manner, where the model’s output logits are masked to only include classes present in the replay buffer.
The outcome of the probe dictates the schedule for the next one. If the model’s performance on the replay buffer exceeds a predefined mastery_threshold, its memory is considered stable and the time until the next probe (replay_gap) is increased by a multiplicative factor. Otherwise, If the performance is below the threshold, the model’s memory is considered weak, and the next probe is scheduled for the very next epoch to encourage more intensive consolidation. The full procedure is detailed below in Algorithm 1.
@@ 2.4.1 Algorithm 1
Require: Raplay Buffer b, Training DataLoader l
Require: Hyperparameters: epochs, mastery_threshold, initial_gap,
gap_multiplier
@@
o.zero_grad()
outputs = m(mixed_data)
loss = c(outputs, mixed_targets)
loss.backward()
o.step()
@@
if epoch == replay_timer and task_id > 0 then
@@
2.5 Code availability
The code for the experiments is available on Github: https://github.com/pritalb/tfc-sr
@@ 3.1 Implementation Details
All experiments were implemented using Pytorch and trained on a single NVIDIA V100 GPU. To avoid randomness and allow fair comparison between all methods, a constant seed (42) was used.
A simple CNN was used for Split MNIST while for the more challenging Split CIFAR, a modified ResNet-18 architecture was used, as detailed in section 2.2. Unless otherwise specified, all models were trained using the Adam optimizer with a learning rate of 0.001, determined via a hyperparameter tuning process on the baseline model. The batch size was 64 for both benchmarks. For Split MNIST, 10 epochs were used per task while 20 were used for Split CIFAR-100.
The key hyperparameters for each continual learning method were selected using coarse hyperparameter tuning to ensure a competitive comparison. For our proposed TFC-SR, we analyzed its sensitivity to both buffer capacity and mastery threshold (see Appendix A and B for full details). As an extension, Buffer size sensitivity was also measured for Standard ER. For the regularization based methods EWC and SI, the regularization strength, $\lambda$, was also selected via a coarse hyperparameter tuning over several orders of magnitude. The final values used for each method are listed in Table 1.
Table 1: Final Hyperparameter values for main experiments
@@ 3.2 Evaluation on Split MNIST
To validate the feasibility of TFC-SR, experiments were conducted on the Split MNIST benchmark. Figure 1 illustrates the learning curves, plotting the average accuracy on all seen tasks after each of the five sequential tasks were completed.
The results clearly demonstrate the severity of catastrophic forgetting. After training on 5 tasks, The baseline (Sequential finetuning) model’s performance dropped to a final accuracy of 10.28%. The regularization methods provided modest improvements over the baseline with EWC scoring 20.37% accuracy while SI got a final accuracy of 32.58%.
A significant performance improvement was observed with the replay-based methods. The standard ER baseline achieved a final accuracy of 90.44%, showing that even non-adaptive replay is highly effective for this benchmark. The proposed TFC-SR method also achieved a strong and comparable final accuracy of 86.22%. Notably, the learning trajectory of TFC-SR was competitive with Standard ER throughout the experiment, briefly outperforming it until the third task before finishing at a similar level by the fourth task. The strong showing of TFC-SR motivated further analysis on a more challenging dataset.
Figure 1: Performance comparison of TFC-SR against baseline methods on the Split MNIST benchmark. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis). Higher is better.
@@ 3.3 Scalability Analysis on Split CIFAR-100
After the feasibility of the proposed TFC-SR was established, It was tested on a much harder Split CIFAR-100 benchmark in order to perform Scalability Analysis. The results are outlined in Figure 2. As expected, all methods experienced a significant drop in performance on this harder benchmark. Sequential fine tuning only managed a final performance of 7.27%. The regularization based methods, EWC and SI, failed to provide any substantial improvement and had final average accuracy of 6.16% and 8.15% respectively. Standard ER, which showed strongest results on previous MNIST benchmark, ended up with a final average accuracy of 7.4%.
@@
Figure 2: Performance comparison of TFC-SR against baseline methods on the Split CIFAR-100 benchmark. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis). Higher is better.
@@ 3.4 Ablation Study: Effect of Buffer Capacity
During the initial hyperparameter tuning for TFC-SR, It was observed that the model’s performance on its memory checks was consistently high, suggesting the memory retention task was not sufficiently challenging with a large buffer. To investigate the behaviour of our method under varying degrees of memory pressure, an ablation study on the replay buffer capacity was conducted. The performance of TFC-SR against Standard ER was compared across four buffer sizes: 100, 500, 1000, and 2000. The full learning curves for each run are detailed in Appendix B, while the final accuracies are summarised in Table 2 and plotted in Figure 3. The results reveal a clear and interesting trade-off between the two methods.
In low-memory (capacity=100) and medium-memory (capacity = 500, 1000) settings, TFC-SR consistently outperforms Standard ER. The performance advantage is more at a capacity of 1000, where TFC-SR achieves a final accuracy of 13.17% against ER’s 7.40%. This suggests that when memory resources are constrained, the active recall probe and adaptive scheduling of TFC-SR provide a significant benefit over a non-adaptive replay strategy.
However, in the high-memory setting (capacity = 2000), this trend reverses. Standard ER achieves a final accuracy of 21.37%, surpassing TFC-SR’s 19.72%. This indicates that when the memory buffer is large and diverse enough, the sheer volume of brute-force replay becomes a highly effective strategy.
@@
Figure 3: Effect of buffer capacity on final model performance. The plot shows the final average accuracy on all tasks seen so far (y-axis) for Standard ER and TFC-SR at different values of buffer capacity (x-axis) on the Split CIFAR-100 benchmark. Higher is better.

### v2

# This manuscript is a preprint and has not been peer reviewed | Version: September 2025

@@ Abstract
Deep neural networks often suffer from a critical limitation known as catastrophic forgetting, where performance on past tasks degrades after learning new ones. This paper introduces a novel continual learning approach inspired by human learning strategies like Active Recall, Deliberate Practice, and Spaced Repetition, named Task-Focused Consolidation with Spaced Recall (TFC-SR). TFC-SR enhances the standard experience replay framework with a mechanism we term the Active Recall Probe. It is a periodic, task-aware evaluation of the model’s memory that stabilizes the representations of past knowledge. We test TFC-SR on the Split MNIST and the Split CIFAR-100 benchmarks against leading regularization-based and replay-based baselines. Our results show that TFC-SR performs significantly better than these methods. For instance, on the Split CIFAR-100, it achieves a final accuracy of 13.17% compared to Standard Experience Replay’s 7.40%. We demonstrate that this advantage comes from the stabilizing effect of the probe itself, and not from the difference in replay volume. Additionally, we analyze the trade-off between memory size and performance and show that while TFC-SR performs better in memory-constrained environments, higher replay volume is still more effective when available memory is abundant. We conclude that TFC-SR is a robust and efficient approach, highlighting the importance of integrating active memory retrieval mechanisms into continual learning systems.
@@ 1 Introduction
Deep neural networks often suffer from a critical limitation known as catastrophic forgetting: performance on previously learned tasks degrades as the model learns to perform newer ones (mccloskeyCohen1989, ). This problem arises from the stability-plasticity dilemma (abraham_memory_2005, ) where a model must be flexible enough to learn newer tasks while simultaneously being stable enough to retain existing knowledge. This seriously affects the ability of a model to perform well in real-world scenarios where the model must adapt to newer data and tasks incrementally over a period of time.
This paper aims to test an approach inspired by effective, synergistic learning strategies employed by humans, such as Active Recall (roediger_test-enhanced_2006, ), Deliberate Practice (ericsson_role_1993, ), and Spaced Repetition (cepeda_spacing_2008, ). Active Recall involves effortfully retrieving information from memory. Deliberate Practice involves studying a task until a certain level of proficiency is achieved. Spaced Repetition refers to reviewing information at increasing intervals with the interval length determined by how well the task has been mastered. For example, if a student is learning Japanese, they may use flashcard applications like Anki to frequently test harder vocabulary (Active Recall), while words that are easier and recalled frequently are studied at progressively longer intervals (Spaced Repetition).
This paper proposes and evaluates a novel methodology to replicate these human learning methods within the continual learning framework. The primary contribution is Task-Focused Consolidation with Spaced Recall (TFC-SR), where the model learns new tasks while concurrently “practicing” previously learned tasks using a form of experience replay. The core of our method is the Active Recall Probe, a periodic "memory check" where the model performs a forward pass on past experiences to evaluate its own memory state. The outcome of this memory check is then used by an adaptive spaced schedule that determines the intensity and frequency of future checks. For this initial investigation, a “naive” notion of mastery is used, where mastery refers to retained performance on previously learned tasks exceeding a predefined threshold. For the purpose of validation, TFC-SR is tested on the Split MNIST and then on the more challenging Split CIFAR-100 benchmarks and compared against standard continual learning baselines, including premier regularization-based methods like EWC (kirkpatrick_overcoming_2017, ) and SI (zenke_continual_2017, ) to demonstrate its effectiveness in alleviating catastrophic forgetting.
@@ 2.1 Task Protocols
To demonstrate the effectiveness of TFC-SR, a simple CNN and a ResNet-18 (he_deep_2016, ) were evaluated on two benchmark datasets, Split MNIST (lecun_gradient-based_1998, ) and Split CIFAR-100 (krizhevsky_learning_2009, ), respectively. In both cases, after the model had learned a task, it had to classify all the classes it had seen so far. This was done to test how well the model retained its old knowledge after new information was made available to it.
Training a CNN on MNIST is a relatively simple task and is often described as the “Hello, World!” of neural networks. A basic implementation can be easily trained to achieve high accuracy on the dataset, making it suitable for testing the feasibility of TFC-SR in a relatively simple setting. MNIST consists of a total of 70,000 grayscale, ${\times}28$ images of handwritten digits 0–9. These 70,000 images are then split into two sets: 60,000 for training and 10,000 for testing. Split MNIST is a variant of this dataset that is further divided into multiple classification tasks. In our case, it was divided into five tasks, each involving the model learning and classifying a pair of digits. This setup was used to demonstrate continual learning, in which the model receives data sequentially rather than having access to the entire training set at once.
CIFAR-100 was chosen to represent a better, relatively more difficult benchmark after TFC-SR demonstrated its effectiveness within a simple setting. CIFAR-100 consists of 60,000 color, ${\times}32$ images representing 100 classes. There are 600 images per class and the 100 classes are grouped into 20 different superclasses. Similar to MNIST, CIFAR-100 is also split into training set and testing set with each containing 50,000 and 10,000 images respectively. Since this was going to be too difficult for a simple CNN to handle, a ResNet-18 was chosen as the model of choice. For our purposes, CIFAR-100 was divided into 10 tasks, with 10 classes per task.
@@ 2.2 Model Architecture
All models were implemented using PyTorch and trained from scratch. For the first benchmark of Split MNIST, a standard CNN was used with convolutional blocks followed by a classifier head. The first convolutional block consisted of a 2D convolutional layer with 32 ${\times}3$ kernels followed by a ReLU activation function and a ${\times}2$ max-pooling layer. The second convolutional block was identical to the first one, with the only difference being that the 2D convolutional layer had 64 ${\times}3$ kernels. The output of these blocks was flattened and passed to the classifier head, which comprised two fully connected layers. The first fully connected layer had 128 units with ReLU activation while the second one was a linear layer with 10 units corresponding to the 10 digit classes.
For the Split CIFAR-100 benchmark, a ResNet-18 model was used. To adapt the model to CIFAR-100’s ${\times}32$ image size, two conventional modifications were made: the initial ${\times}7$ convolutional layer was replaced with a ${\times}3$ convolutional layer, and the subsequent max-pooling layer was removed by replacing it with an identity mapping. The final fully connected layer was replaced with a linear layer containing 100 units, corresponding to the 100 classes in the CIFAR-100 dataset.
@@ 2.3 Learning Algorithms and Baselines
  • Sequential Fine-tuning (Baseline): This baseline represents a lower-bound reference without any additional method and was used to evaluate the benefit of applying TFC-SR.
  • Standard Experience Replay (ER) (mnih_playing_2015, ): In this method, the model has access to data from previous tasks. Since TFC-SR is also a replay-based approach, Standard ER serves as a reference for assessing the contribution of the Active Recall Probe.
  • Elastic Weight Consolidation (EWC) (kirkpatrick_overcoming_2017, ): A leading regularization-based method that adds a penalty term to protect weights that are important for previous tasks, as determined by the Fisher Information Matrix.
  • Synaptic Intelligence (SI) (zenke_continual_2017, ): This is another prominent regularization-based method. SI protects weights that have contributed significantly to reducing the loss during training on prior tasks. The regularization strength, $\lambda$, was determined via coarse hyperparameter tuning for both EWC and SI.
@@ 2.4 TFC-SR algorithm
TFC-SR extends the standard experience replay framework using an adaptive scheduling mechanism based on Active Recall and Spaced Repetition. The training process for each task consists of two main components: continuous mixed-batch training and an adaptive active recall schedule.
Continuous Mixed-Batch Training: for all tasks, the model is trained on mixed batches of data. Each batch is composed of a 50/50 split of new examples from the current task and examples randomly sampled from a replay buffer containing data from all the past tasks. To maintain a diverse and representative memory of the past experiences within a fixed memory budget, a reservoir replay buffer of fixed capacity was used.
Active Recall Probe and Adaptive Scheduling: The core novelty of TFC-SR is its adaptive scheduling mechanism, which we term the Active Recall Probe. This probe is a “memory check” performed periodically after each training epoch. It involves evaluating the model’s performance on a batch of samples drawn from the replay buffer. To obtain an accurate measure of retained knowledge in a class-incremental setting, this evaluation is conducted in a task-aware manner, where the model’s output logits are masked to include only the classes currently present in the replay buffer. We refer to this evaluation metric as the “average accuracy” throughout the paper, indicating the model’s overall accuracy on the combined set of all seen tasks, rather than an arithmetic mean of accuracies computed separately per task.
The outcome of the probe dictates the schedule for the next one. If the model’s performance on the replay buffer exceeds a predefined mastery_threshold, its memory is deemed stable and the time until the next probe (replay_gap) is increased by a multiplicative factor. Otherwise, if the performance is below the threshold, the model’s memory is considered weak and the next probe is scheduled for the very next epoch to encourage more intensive consolidation. The full procedure for training on a single task (i.e., one experience) is detailed below in Algorithm 1.
@@ 2.4.1 Algorithm 1
Require: Replay Buffer b
Require: Hyperparameters:
epochs, mastery_threshold, initial_gap, gap_multiplier
@@
// here, sufficient means there is enough data for
// a 50/50 mixed batch
@@
// TrainStep: performs one gradient update using
// the given inputs and targets
TrainStep(m, o, c, mixed_data, mixed_targets)
@@
if epoch == replay_timer and b not empty then
// perform an active recall probe
@@ 3.1 Implementation Details
All experiments were implemented using PyTorch and trained on a single NVIDIA V100 GPU via Google Colab. To reduce randomness and ensure fair comparisons, a fixed random seed (42) was used.
A simple CNN was used for Split MNIST, while for the more challenging Split CIFAR-100, a modified ResNet-18 architecture was used, as described in Section 2.2. Unless otherwise specified, all models were trained using the Adam optimizer with a learning rate of 0.001, determined via a hyperparameter tuning process on the baseline model. The batch size was set to 64 for both benchmarks. For Split MNIST, 10 epochs were used per task, while 20 were used for Split CIFAR-100.
The key hyperparameters for each continual learning method were selected using coarse hyperparameter tuning to ensure a competitive comparison. For our proposed TFC-SR, we analyzed its sensitivity to both the mastery threshold and the buffer capacity (see Appendix A and B for full details). As an extension, buffer size sensitivity was also measured for Standard ER. For the regularization-based methods EWC and SI, the regularization strength, $\lambda$, was also selected via coarse hyperparameter tuning over several orders of magnitude. The final values used for each method are listed in Table 1.
Table 1: Final hyperparameter values for main experiments.
@@ 3.2 Evaluation on Split MNIST
To validate the feasibility of TFC-SR, experiments were conducted on the Split MNIST benchmark. Figure 1 illustrates the learning curves, plotting the average accuracy on all seen tasks after each of the five sequential tasks was completed.
The results clearly demonstrate the severity of catastrophic forgetting. After training on five tasks, the performance of the baseline (Sequential Fine-tuning) model dropped to a final accuracy of 10.28%. The regularization-based methods provided modest improvements over the baseline, with EWC achieving 20.37% accuracy and SI reaching a final accuracy of 32.58%.
A significant performance improvement was observed with the replay-based methods. The standard ER baseline achieved a final accuracy of 90.44%, showing that even non-adaptive replay is highly effective for this benchmark. The proposed TFC-SR method also achieved a strong and comparable final accuracy of 86.22%. Notably, the learning trajectory of TFC-SR was competitive with Standard ER throughout the experiment, briefly outperforming it until the third task before finishing at a similar level by the fourth task. The strong performance of TFC-SR motivated further analysis on a more challenging dataset.
Figure 1: Performance comparison of TFC-SR against baseline methods on the Split MNIST benchmark. The plot shows average accuracy on all previously seen tasks (y-axis) after each new task is learned (x-axis). Higher is better.
@@ 3.3 Scalability Analysis on Split CIFAR-100
After the feasibility of the proposed TFC-SR was established, it was tested on a much harder Split CIFAR-100 benchmark in order to perform Scalability Analysis. The results are outlined in Figure 2. As expected, all methods experienced a significant drop in performance on this harder benchmark. Sequential Fine-tuning only managed a final performance of 7.27%. The regularization-based methods, EWC and SI, failed to provide any substantial improvement and had final average accuracy of 6.16% and 8.15% respectively. Standard ER, which showed strongest results on previous MNIST benchmark, ended up with a final average accuracy of 7.4%.
@@
Figure 2: Performance comparison of TFC-SR against baseline methods on the Split CIFAR-100 benchmark. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learned (x-axis). Higher is better.
@@ 3.4 Ablation Study: Effect of Buffer Capacity
During the initial hyperparameter tuning for TFC-SR on the Split CIFAR-100 benchmark, it was observed that the model’s performance on memory checks was consistently high, suggesting that the memory retention task was not sufficiently challenging when using a relatively large buffer. To investigate the behavior of our method under varying degrees of memory pressure, an ablation study was conducted on the replay buffer capacity. The performance of TFC-SR against Standard ER was compared across four buffer sizes: 100, 500, 1000, and 2000. The complete learning curves for each run are detailed in Appendix B, while the final accuracies are summarized in Table 2 and visualized as a trend plot in Figure 3, with the buffer capacity shown on a logarithmic scale. The results reveal a clear and meaningful trade-off between the two methods.
In low-memory (capacity = 100) and medium-memory (capacity = 500, 1000) settings, TFC-SR consistently outperforms Standard ER. The performance advantage is greater at a capacity of 1000, where TFC-SR achieves a final accuracy of 13.17% against Standard ER’s 7.40%. This suggests that when memory resources are constrained, the Active Recall Probe and adaptive scheduling of TFC-SR provide a significant benefit over a non-adaptive replay strategy.
However, in the high-memory setting (capacity = 2000), this trend reverses. Standard ER achieves a final accuracy of 21.37%, surpassing TFC-SR’s 19.72%. This outcome suggests that in high-memory settings, the diversity of the buffer alone is sufficient to maintain performance, reducing the relative advantage of TFC-SR’s Active Recall Probe and adaptive scheduling mechanisms.
@@
Figure 3: Effect of replay buffer capacity on final average accuracy. The plot shows the final average accuracy (y-axis) after training on 10 tasks for Standard ER and TFC-SR across varying buffer capacities (x-axis, log-scaled base 10) on the Split CIFAR-100 benchmark. Higher values indicate better performance.